Katie's environmental club has bolunteered to clean up a section of highway. They can clean up to about one mile of highway in 6 hours. If they work 8 hours a day, how many days will it take them to clean up 12 miles of highway
Answer by asuar010(324)

we set up proportions and we get 1/6=12/x so x = 48 hrs so it will take them 2 days